OpenStudy (anonymous):

can someone help me with a couple of problems. I just started this class and i am lost. For the given equation, list the intercepts and test for Symmetry. y=x^3-1000

OpenStudy (fellowroot):

for x int set y=0 for y int set x=0, symmetry is when you look at the graph and see if its reflects over the y axis or x axis or orgin

OpenStudy (anonymous):

this equation is not symmetric about any axes

OpenStudy (anonymous):

\( y=x^3-1000=(x-10)(x^2+10x+100)\) only zero is at \(x=10\)